# Understanding Dataset Distillation via Spectral Filtering

### v1

@@ 4.4 Ablation Studies
We conducted two experiments to demonstrate the importance of each module in the proposed model. The first experiment, shown in Table 5, is used to identify the effectiveness of different loss functions. Specifically, we sequentially remove the classification loss, i.e., $\mathcal{L}_{\text{cls}}$ and signal-matching loss, i.e., $\mathcal{L}_{\text{signal}}$, and evaluate CFM on three datasets, including CIFAR-100, Tiny-ImageNet, and ImageNet-1k. We can observe that all three loss functions contribute to the performance of CFM, and the filter-matching loss plays a fundamental role. This discovery suggests that the important knowledge of the real datasets may encoded in the FFC matrix of the feature maps.
The second experiment is used to verify the effectiveness of the curriculum strategy. Specifically, we evaluate the performance of different filters, including low-pass filter, high-pass filter with parameter $\beta=1e^{-1},1e^{-2},1e^{-3},1e^{-4}$, and high-pass filter with CFM. The results are shown in Figure 3, from which we can find that the performance of the low-pass filter is far away from the high-pass filters, indicating that the high-frequency information is more important for DD. However, the extremely high frequency will also lead to performance degradation. For example, the performance of $\beta=1e^{-4}$ is not as good as other parameters. On the other hand, CFM balances the proportion between low-frequency and high-frequency information by setting a proper frequency band, consistently outperforming other filters.
@@
4.5 Visualization
To provide a more intuitive comparison of various DD methods, we visualize the synthetic images generated by SRe2L, G-VBSM, and CFM in Figure 4. It can be observed that SRe2L and G-VBSM struggle to produce meaningful images, particularly with the low-resolution Tiny-ImageNet dataset. In contrast, the images synthesized by CFM exhibit clear semantic information, e.g., recognizable shapes, demonstrating the effectiveness of matching frequency-specific features between the real and synthetic datasets.
@@ 5 Related Work
As we have analyzed the objectives of DD in Section 2, we only introduce some orthogonal research lines, including data parameterization, model augmentation, and theory.
Data Parameterization.
Traditionally, the synthetic images are optimized in the pixel space, which is inefficient due to the potential data redundancy. Therefore, some methods explore how to parameterize the synthetic images. IDC [IDC] distillates low-resolution images to save budget and up-samples them in the evaluation stage. Haba [Haba] designs a decoder network to combine different latent codes for diverse synthetic datasets. RDED [RDED] stitches the core patches of real images together as the synthetic data. Some methods use generative models to decode the synthetic datasets from latent codes, including GAN-based methods [IT-GAN, MGDD] and diffusion-based methods [diff1, diff2]. Generally, data parameterization can be incorporated with different DD objectives to reduce their storage budgets and improve efficiency.
Model Augmentation.
The knowledge of the datasets is mostly encoded in the pre-train distillation networks. Therefore, some methods focus on training a suitable network for DD by designing some augmentation strategies. FTD [FTD] constrains model weights to achieve a flat trajectory and reduces the accumulated errors in MTT. model_aug uses early-stage models and parameter perturbation to increase the search space of the parameters. More recently, DWA [DWA] employs directed weight perturbations on the pre-train models, which not only maintains the unique features of each synthetic data but also improves the diversity.
Theory.
Some works tend to explore DD from a theoretical perspective. For example, theorem_size analyzes the size and approximation error of the synthetic datasets. theorem_bias explores the influence of spurious correlations in DD. theorem_learning explains the information captured by the synthetic datasets. theorem_core provides a formal definition of DD and provides a foundation analysis on the optimization of DD. Although these methods have some important insights, they fail to explain the roles of existing methods. In contrast, UniDD establishes the relationship between DD objectives and spectral filtering.
@@ 6 Conclusion
In this paper, we introduce UniDD, a framework that unifies various DD objectives through spectral filtering. UniDD demonstrates that each DD objective corresponds to a filter function applied on the FFC and FLC matrices. This finding reveals the nature of various DD methods and inspires the design of new methods. Based on UniDD, we propose CFM to encode both low- and high-frequency information by gradually changing the filter parameter. Experiments conducted on various datasets validate the effectiveness and generalization of the proposed method. A promising future direction is to generalize UniDD to the distillation of unsupervised and multi-modal datasets.
Supplementary Material
1 Derivation of UniDD
@@
2 Implementation Details.
2.1 Experimental Setup
We use the squeeze, recover, and relabel pipeline, introduced by SRe2L, for the distillation of CFM. The details of datasets, setup (e.g., optimizer), and hyper-parameters are listed in Tables 1, 2, 3, and 4.
@@ 2.1 Experimental Setup
Resolution 64$\times$64 64$\times$64 64$\times$64 224$\times$224
@@
3 Visualization
Finally, we visualize the synthetic images of ImageNet-1k to provide a comprehensive exhibition.
Figure 5: Synthetic images of ImageNet-1k. From left to right, the frequency gradually increases.

### v2

@@ 6.4 Ablation Studies
We conducted two experiments to demonstrate the importance of each module in the proposed model. The first experiment, shown in Table 6, is used to identify the effectiveness of different loss functions. Specifically, we sequentially remove the classification loss and signal-matching loss, and evaluate CFM on three datasets, including CIFAR-100, Tiny-ImageNet, and ImageNet-1k. We can observe that all three loss functions contribute to the performance of CFM, and the filter-matching loss plays a fundamental role. This discovery suggests that important knowledge of the real datasets may be encoded in the FFC matrix of the feature maps.
The second experiment is used to verify the effectiveness of the curriculum strategy. Specifically, we evaluate the performance of different filters, including low-pass filter, high-pass filter with parameter $\beta=\{1e^{-1},e^{-2},1e^{-3},1e^{-4}\}$, and high-pass filter with CFM. The results are shown in Figure 6, from which we can find that the performance of the low-pass filter is far away from the high-pass filters, indicating that the high-frequency information is more important for DD. However, the extremely high frequency will also lead to performance degradation. For example, the performance of $\beta=1e^{-4}$ is not as good as other parameters. On the other hand, CFM balances the ratio between low-frequency and high-frequency by setting a proper frequency band, consistently outperforming other filters.
6.5 Time and Space Overhead
We compare the time and space overhead of CFM with two representative DD methods: SRe2L SRe2L and G-VBSM G-VBSM in the ImageNet-1k dataset with a ResNet‑18 backbone. To ensure fairness, all methods optimize 500 synthetic images per step. The results are averaged over 1,000 iterations on a RTX 4090 GPU. Results are shown in Tables 8 and 8.
All methods use the basic classification loss. SRe2L proposes a batch normalization loss, $\mathcal{L}_{\text{BN}}$, to align the mean and variance statistics. Building on this, G-VBSM further adds a densification loss, $\mathcal{L}_{\text{dense}}$, to improve the diversity of the synthesized images. Unlike them, CFM uses the filter-matching and signal-matching losses. Compared to SRe2L, CFM has a slightly higher time overhead, but it is still lower than G-VBSM. On the other hand, CFM has the lowest space overhead, as the number of convolutional layers is less than that of batch normalization, demonstrating the efficiency of CFM.
6.6 Visualization
@@
To provide a more intuitive comparison of various DD methods, we visualize the synthetic images generated by SRe2L, G-VBSM, and CFM in Figure 5. It can be observed that SRe2L and G-VBSM struggle to produce meaningful images, particularly with the low-resolution Tiny-ImageNet dataset. In contrast, the images synthesized by CFM exhibit clear semantic information, e.g., recognizable shapes, demonstrating the effectiveness of matching frequency-specific features between the real and synthetic datasets. See Appendix D for more visualizations.
@@ 7 Related Work
Data Parameterization. Traditionally, the synthetic images are optimized in the pixel space, which is inefficient due to the potential data redundancy. Therefore, some methods explore how to parameterize the synthetic images. IDC (IDC) distillates low-resolution images to save budget and up-samples them in the evaluation stage. Haba (Haba) designs a decoder network to combine different latent codes for diverse synthetic datasets. RDED (RDED) stitches the core patches of real images together as the synthetic data. Some methods use generative models to decode the synthetic datasets from latent codes, including GAN-based methods (IT-GAN; MGDD), diffusion-based methods (diff1; diff2), and implicit function-based methods (FreD). Generally, data parameterization can be incorporated with different DD objectives to improve their efficiency.
Model Augmentation. The knowledge of the datasets is mostly encoded in the pre-trained distillation networks. Therefore, some methods focus on training a suitable network for DD by designing some augmentation strategies. FTD (FTD) constrains model weights to achieve a flat trajectory and reduces the accumulated errors. model_aug use early-stage models and parameter perturbation to increase the search space of the parameters. More recently, DWA (DWA) employs directed weight perturbations on the pre-training model that maintain the unique features of each synthetic data.
Theory. Some works tend to explore DD from a theoretical perspective. For example, theorem_size analyzes the size and approximation error of the synthetic datasets. theorem_bias explores the influence of spurious correlations in DD. theorem_learning explains the information captured by the synthetic datasets. theorem_core provides a formal definition of DD and gives a foundation analysis on the optimization of DD. These methods have some important insights, but fail to explain the roles of existing methods. In contrast, UniDD establishes the relationship between DD objectives and spectral filtering.
@@ 8 Conclusion
In this paper, we introduce UniDD, a framework that unifies various DD objectives through spectral filtering. UniDD demonstrates that each DD objective corresponds to a filter function applied to the FFC and FLC matrices. This finding reveals the nature of various DD methods and inspires the design of new methods. Based on UniDD, we propose CFM to encode both low- and high-frequency information by gradually changing the filter parameter. Experiments conducted on various datasets validate the effectiveness and generalization of the proposed method. A promising future direction is to generalize UniDD to the distillation of unsupervised and multi-modal datasets.
ACKNOWLEDGMENT
This project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Award Number: MOE-T2EP20122-0006).
ETHICS STATEMENT
This work aims to provide a unified view of dataset distillation and deepen our understanding of data efficiency. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
REPRODUCIBILITY
To ensure reproducibility, we provide additional implementation details in Appendix C. During the review process, we submit the source code as supplementary material, and we promise to release the code publicly upon acceptance of the paper.
Appendix A Statement on LLM Usage
In drafting this manuscript, LLMs were used exclusively for polishing language and checking grammar. The content was not directly copied from LLM outputs, and all scientific concepts, analyses, and conclusions represent the authors’ original work.
Appendix B Derivation of UniDD
@@
Appendix C Implementation Details.
C.1 Algorithm
Algorithm 1 illustrates the distillation process of CFM.
Algorithm 1 Curriculum Frequency Matching
2:Synthetic dataset $\mathcal{S}=(H_{s},Y_{s})$
3:Feed $H$ into $\phi$ for a forward pass
4:for layer index $l=1,\cdots,L$ do
5:  Calculate $\Psi^{l}$ and $\Phi^{l}$ based on Eq. 16
6:end for
7:for batch index $b=1,\cdots,B$ do
8:  Initialize $H_{s}^{b}$ with randomly sampled real data
9:  Update $\beta_{b}$ based on Eq. 18
10:  repeat
11:   Feed $H_{s}^{b}$ into $\phi$ for a forward pass
12:   for layer index $l=1,\cdots,L$ do
13:     Calculate $\Psi_{s}^{l,b}$ and $\Phi_{s}^{l,b}$ based on Eq. 17
15:   end for
16:   Calculate $\mathcal{L}_{\text{ce}}$ and back-propagate $\mathcal{L}$
17:   Update $H_{s}^{b}$
18:  until Reached the number of iteration $\mathcal{I}$
19:end for
C.2 Environment
All experiments are conducted on a single GeForce RTX 4090.
C.3 Experimental Setup
We use the squeeze, recover, and relabel pipeline, introduced by SRe2L, for the distillation of CFM. The details of datasets, setup (e.g., optimizer), and hyper-parameters are listed in Tables 9, 10, 11, and 12.
@@ C.3 Experimental Setup
Resolution $\times64$ $\times64$ $\times64$ $\times224$
@@
Appendix D Visualization
Finally, we visualize more synthetic images of ImageNet-1k in Figure 6 to provide a comprehensive exhibition. Specifically, each column represents the images sampled from the same batch. From left to right, the value of $\beta_{d}$ decreases, and the high-frequency information gradually increases. It can be observed that images on the left side are somewhat blurry, while images on the right side have more complex textures, validating the conclusions in Section 4.
Figure 6: Synthetic images. From left to right, the high-frequency gradually increases.